# Genomic Next-Token Predictors are In-Context Learners

### v2

@@ 1 Introduction
Scaling Large Language Models (LLMs) has revealed an unexpected and powerful capacity: in-context learning (ICL) (radford2018improving; brown2020language), the ability to infer and apply abstract patterns purely from examples contained within their input. Unlike conventional adaptation, which relies on gradient updates, LLMs internalize mechanisms for flexible pattern induction (olsson2022context), allowing them to perform few-shot generalization and analogical reasoning (srivastava2023beyond), without explicit parameter updates. This behavior, which arises organically from large-scale “next-token” prediction on human language, has been interpreted as emergent forms of meta-learning (learning to learn) within a fixed parametric system.
Almost all prominent evidence of emergent ICL so far (srivastava2023beyond, inter alia) comes from training on human language (e.g., English). This raises a fundamental question: is there something inherently special about human language that enables ICL to emerge? One can take two different stances on this:
  • H1: Human language possesses distinctive distributional properties such as parallelism or compositionality (chen2024parallel; hahn2023theory) that uniquely nurture ICL.
@@
While many works have examined H1, the veracity of H2 remains far less understood. To this end, we turn to a radically different substrate: the genome. Genomic sequences can be viewed as a form of natural language that has evolved through nature. Like human language, they comprise complex symbol sequences that exhibit rich statistical regularities – motifs, repeats, dependencies (benegas2025genomic) that could, in principle, support pattern induction within context.
Testing for the existence of ICL requires large-scale models. Historically, genomic models were small and specialized, optimized for narrow biological tasks such as motif discovery or structure prediction. The advent of deep biological models (Ji2020DNABERTPB) has expanded their scale, paralleling those of early LLMs. Most recently, the Evo2 series (evo2report) marks a major step forward: a large-scale model trained solely on ‘next-nucleotide’ prediction. With training scale comparable to mid-sized LLMs (e.g., Qwen3-14B (yang2025qwen3technicalreport)), these genomic models finally make it possible to systematically compare genomic and linguistic representations under large-scale autoregressive training. If ICL arises primarily from scale and predictive compression (H2), then large-scale genomic models such as Evo2 must already exhibit it under our noses. Like LLMs, they are trained for next-token prediction without any explicit meta-learning signal. Yet, despite extensive analysis of Evo2’s biological capabilities (e.g., protein fitness prediction, variant effect estimation), its potential for emergent few-shot or analogical reasoning within context remains unexplored.
@@
We further perform analyses to compare the types of tasks at which genomic and linguistic models excel. Our results reveal distinct patterns: genomic models learn faster with more demonstrations, while linguistic models gain more from scale (§4.2); genomic models perform best on direct copying or local transformations, whereas linguistic models excel on tasks requiring single-bit dependencies or global summary statistics (§4.3, §4.4). We note the caveat that these findings are specific to our experimental conditions (e.g., our choice of reasoning tasks) and may not be a holistic statement about ICL in genomic and linguistic models.
@@ 2 Related Work and Broader Context
Why does ICL remain interesting? ICL remains a striking and conceptually rich phenomenon: it emerges without task-specific training yet enables rapid, in-situ adaptation. It comes closest to long-standing ambitions in classical AI—such as case-based (aamodt1994case; lancaster1987problem; ross1984remindings; ellman1989explanation) and analogical reasoning (hofstadter2001analogy; gentner2017analogy; holyoak2001place) where agents solve new problems by reusing solutions from prior (analogous) examples.
@@
Beyond its theoretical appeal, ICL has become a practical workhorse in recent advances, e.g., synthetic data generation using pretrained models (wang2023selfinstruct; shao2023synthetic; tunstall2023zephyr) or reasoning (wei2022chain; nye2021show; yao2022react; yasunaga2023large). In short, ICL is both a window into emergent learning dynamics and a cornerstone for building adaptive, interpretable systems.
@@
Emergent-ICL $\neq$ Meta-ICL: A parallel line of research studies models trained with an explicit ICL objective – learning to predict over collections of input–output pairs $(x,f(x))$. For example, regression models that are trained explicitly to infer an underlying rule (e.g., linear, polynomial, or sinusoidal) from few-shot examples (garg2022can; li2023transformers; raventos2024pretraining; nejjar2024context). This setup mirrors meta-learning, where a model is explicitly “trained to learn” (finn2017model; bertinetto2018meta; zintgraf2019fast). We refer to such systems as Meta-ICL (panwar2023context; min2022metaicl; wu2022continued; kirsch2022general; zhang2023incontext), in contrast to emergent ICL in large pretrained models, which arises organically from next-token prediction over natural data.
@@
Meta-ICL results demonstrate that transformers can be optimized to perform meta-learning but fail to explain why emergent ICL appears in models never trained for it. Critically, Meta-ICL generalizes poorly outside its training domain (naim2024re). For example, a Meta-ICL model trained on linear and cosine functions fails on their composition (yadlowsky2023pretraining)—and does not capture higher-order abilities like in-context reasoning (wei2022chain; kojima2022large) or adaptive self-refinement (madaan2023self) – unless explicitly trained for them. In contrast, emergent ICL in LMs generalizes broadly and flexibly across diverse problem types (srivastava2023beyond). These behavioral and mechanistic differences support our working assumption that Meta-ICL and emergent ICL are fundamentally distinct phenomena (shen2024icl_vs_gd; mishra2025ia2alignmenticlactivations). Throughout, our focus in this work is on the latter – the organic emergence of in-context learning in large pretrained models.
@@
Are there any prior results on emergent ICL in genomic models? Most genomic models to date are encoders (Ji2020DNABERTPB; Dallatorre2024NucleotideTB), which preclude autoregressive generation and the study of ICL in the standard sense. Among the few autoregressive genomic models, most remain small, e.g., GENA-LM (300M) (Fishman2024GENALMAF) and scGPT (53M) (cui2024scgpt). The recently-released Evo2 series (evo2report) includes models up to 40B parameters, comparable in scale to modern LLMs. This model family provides the first viable testbed for probing emergent ICL, which we study here. A related effort, HyenaDNA (nguyen2023hyenadna), trains on million-token genomic sequences and reports ICL-like behaviors; however, these involve meta-ICL mechanisms (e.g., soft prompting and instruction fine-tuning; see their Sec. 4.3), rather than organically emergent ICL. An additional avenue of research demonstrates that genomic models can perform a form of ICL by copying prior tokens in-context for prediction (kantroo2025incontextlearningdistortrelationship) – akin to how induction heads function (olsson2022incontextlearninginductionheads). This, however, doesn’t demonstrate any ability to infer latent functions or apply a novel rule in-context.
@@
Are there instances of emergent ICL in other non-linguistic modalities? While transformers have been extended to many modalities beyond language, genuine cases of emergent ICL arising purely from large-scale pretraining remain rare. In neuro-signal modeling, for instance, EEG-GPT (kim2024eeg) is trained with explicit in-context demonstrations, making it an instance of Meta-ICL rather than emergent adaptation. Similarly, multimodal models such as Flamingo (alayrac2022flamingo), BLIP/BLIP-2 (li2022blip; li2023blip2), and Emu (sun2023generative; sun2024generative) exhibit in-context behavior only because they are explicitly trained to process demonstration patterns – again, meta-learned rather than emergent. To our knowledge, the only claimed instance of emergent ICL outside language is in vision:  bai2024sequential show that transformers pretrained on natural visual sequences can infer analogical and compositional tasks without explicit supervision. The weakness of this result is that their ICL prompts consist of long sequences of contiguous frames (bai2024sequential, Fig 10), which effectively reduces the task to next-frame prediction, rather than true ICL that requires inferring compositions of multiple demonstrations.
@@ 3.1 Experimental Desiderata
Desideratum 1: Cross-domain comparability. Our experimental framework requires that each task be performable by both language and genomic models. Thus, every task must be representable in both linguistic and nucleotide alphabets. This constraint excludes existing benchmarks that rely on domain-specific semantics – such as language reasoning datasets (e.g., BIG-Bench, StrategyQA, GSM8K (srivastava2023beyond; mor2021strategyqa; cobbe2021gsm8k)) or biological tasks (e.g. variant effect prediction, exon identification  (evo2report)). While one could, in principle, translate domain-specific tasks into an alternate alphabet (e.g., mapping language tokens to base sequences via quaternary encoding), doing so inherently biases the evaluation: the source domain retains advantages, while the target domain must operate on representations that are unnatural to it. As a result, such cross-domain encodings confound the comparison, reflecting representational translation artifacts rather than genuine differences in ICL behavior.
Desideratum 2: Limited vocabulary. Since genomic models operate on only four nucleotides (A, T, C, G), tasks must be expressible within an equally compact alphabet. This rules out the existing symbolic reasoning and analogy benchmarks (hodel2023response; lewis2024evaluating; webb2025evidence), which rely on richer vocabularies (e.g., shapes, colors, or linguistic tokens with explicit semantic roles). Such tasks cannot be faithfully represented in a four-token regime without introducing artifacts or structural loss. Accordingly, our evaluation focuses on tasks that retain abstract reasoning structure while remaining compatible with the low-vocabulary symbolic space shared across linguistic and genomic models.
@@ 3.2 Experimental Setup
Task Formulation: in-context program induction over abstract symbols. Given our desiderata (§3.1), we design symbolic induction tasks that require a small lexicon. Each task requires a model to infer a hidden transformation rule from a few input–output examples and apply it to a new query. We refer to this process as program synthesis in-context. This setup isolates in-context learning ability by removing the influence of pretrained world knowledge. If the function family $f$ and symbol space $S$ are chosen carefully, the evaluation is far from the model’s pretraining distribution. Similar formulations have been used in prior work on compositional reasoning (gpt3fewshot), analogical reasoning via Raven’s Progressive Matrices (Raven1962ManualFR; webb2023emergent), and more recently in ARC-AGI (chollet2019measure).
@@ 3.3 Model Families and Selection Rationale
For fair cross-domain comparison, we use two representative model families: Qwen3 for human language and Evo2 for genomics. The rationale for this selection is as follows: (a) Parameter scaling: Both families span multiple orders of magnitude in parameter count, enabling systematic scaling analysis of ICL ability. The Qwen3 series ranges from 0.6B to 14B parameters, while Evo2 includes 1B, 7B, and 40B models (yang2025qwen3technicalreport; evo2report). This parallel scaling structure facilitates consistent measurement of how ICL performance evolves with model size across linguistic and genomic modalities. (b) Compute matching: The largest models in each family are trained with comparable total compute, offering an opportunity for an approximately compute-matched cross-domain comparison. Using the standard $6ND$ estimate (kaplan2020scaling), Qwen3-14B-Base is trained with about $3.2\times 10^{24}$ FLOPs, while Evo2-40B is trained with $2.25\times 10^{24}$ FLOPs (yang2025qwen3technicalreport; evo2report), making Evo2 uniquely well-suited for comparison with Qwen3 at scale. (c) Availability of base models: The Qwen3 family releases base (pre-instruction-tuned) models at all scales up to 14B parameters, enabling direct evaluation of the intrinsic inductive reasoning ability of pure next-token predictors, without instruction-tuning artifacts. The Evo2 base models have no additional instruction tuning applied – the only nuance is that Evo2 1B is trained at a context length of 8192 nucleotides, whereas the 7B/40B are extended to a context length of one million. (d) Licensing: All models are released under Apache 2.0 (yang2025qwen3technicalreport; arcinstitute_evo2_1b_base_hf_2025; arcinstitute_evo2_7b_hf_2025; arcinstitute_evo2_40b_hf_2025), ensuring reproducibility.

### v3

@@ 1 Introduction
Scaling Large Language Models (LLMs) has revealed an unexpected and powerful capacity: in-context learning (ICL) (Radford et al., 2018; Brown et al., 2020b), the ability to infer and apply abstract patterns purely from examples contained within their input. Unlike conventional adaptation, which relies on gradient updates, LLMs internalize mechanisms for flexible pattern induction (Olsson et al., 2022b), allowing them to perform few-shot generalization and analogical reasoning (Srivastava et al., 2023), without explicit parameter updates. This behavior, which arises organically from large-scale “next-token” prediction on human language, has been interpreted as emergent forms of meta-learning (learning to learn) within a fixed parametric system.
Almost all prominent evidence of emergent ICL so far (Srivastava et al., 2023, inter alia) comes from training on human language (e.g., English). This raises a fundamental question: is there something inherently special about human language that enables ICL to emerge? One can take two different stances on this:
  • H1: Human language possesses distinctive distributional properties such as parallelism or compositionality (Chen et al., 2024; Hahn and Goyal, 2023) that uniquely nurture ICL.
@@
While many works have examined H1, the veracity of H2 remains far less understood. To this end, we turn to a radically different substrate: the genome. Genomic sequences can be viewed as a form of natural language that has evolved through nature. Like human language, they comprise complex symbol sequences that exhibit rich statistical regularities – motifs, repeats, dependencies (Benegas et al., 2025) that could, in principle, support pattern induction within context.
Testing for the existence of ICL requires large-scale models. Historically, genomic models were small and specialized, optimized for narrow biological tasks such as motif discovery or structure prediction. The advent of deep biological models (Ji et al., 2021) has expanded their scale, paralleling those of early LLMs. Most recently, the Evo2 series (Brixi et al., 2025) marks a major step forward: a large-scale model trained solely on ‘next-nucleotide’ prediction. With training scale comparable to mid-sized LLMs (e.g., Qwen3-14B (Yang et al., 2025)), these genomic models finally make it possible to systematically compare genomic and linguistic representations under large-scale autoregressive training. If ICL arises primarily from scale and predictive compression (H2), then large-scale genomic models such as Evo2 must already exhibit it under our noses. Like LLMs, they are trained for next-token prediction without any explicit meta-learning signal. Yet, despite extensive analysis of Evo2’s biological capabilities (e.g., protein fitness prediction, variant effect estimation), its potential for emergent few-shot or analogical reasoning within context remains unexplored.
@@
We further perform analyses to compare the types of tasks at which genomic and linguistic models excel. Our results reveal distinct patterns: genomic models learn faster with more demonstrations, while linguistic models gain more from scale (§4.2); genomic models perform best on direct copying or local transformations, whereas linguistic models excel on tasks requiring single-bit dependencies or global summary statistics (§4.3, §4.4). We note the caveat that these findings are specific to our experimental conditions (e.g., our choice of reasoning tasks) and may not be a holistic statement about ICL in genomic and linguistic models. We further validate these findings by demonstrating that Evo2 also exhibits robust few-shot learning on a real-world genomic promoter classification task (Appendix E).
@@ 2 Related Work and Broader Context
Why does ICL remain interesting? ICL remains a striking and conceptually rich phenomenon: it emerges without task-specific training yet enables rapid, in-situ adaptation. It comes closest to long-standing ambitions in classical AI—such as case-based (Aamodt and Plaza, 1994; Lancaster and Kolodner, 1987; Ross, 1984; Ellman, 1989) and analogical reasoning (Hofstadter, 2001; Gentner and Hoyos, 2017; Holyoak et al., 2001) where agents solve new problems by reusing solutions from prior (analogous) examples.
@@
Beyond its theoretical appeal, ICL has become a practical workhorse in recent advances, e.g., synthetic data generation using pretrained models (Wang et al., 2023c; Shao et al., 2023; Tunstall et al., 2023) or reasoning (Wei et al., 2022; Nye et al., 2021; Yao et al., 2023; Yasunaga et al., 2023). In short, ICL is both a window into emergent learning dynamics and a cornerstone for building adaptive, interpretable systems.
@@
Emergent-ICL $\neq$ Meta-ICL: A parallel line of research studies models trained with an explicit ICL objective – learning to predict over collections of input–output pairs $(x,f(x))$. For example, regression models that are trained explicitly to infer an underlying rule (e.g., linear, polynomial, or sinusoidal) from few-shot examples (Garg et al., 2022; Li et al., 2023c; Raventós et al., 2023; Nejjar et al., 2024). This setup mirrors meta-learning, where a model is explicitly “trained to learn” (Finn et al., 2017; Bertinetto et al., 2019; Zintgraf et al., 2019). We refer to such systems as Meta-ICL (Panwar et al., 2023; Min et al., 2022; Wu et al., 2022; Kirsch et al., 2022; Zhang et al., 2023), in contrast to emergent ICL in large pretrained models, which arises organically from next-token prediction over natural data.
@@
Meta-ICL results demonstrate that transformers can be optimized to perform meta-learning but fail to explain why emergent ICL appears in models never trained for it. Critically, Meta-ICL generalizes poorly outside its training domain (Naim et al., 2024). For example, a Meta-ICL model trained on linear and cosine functions fails on their composition (Yadlowsky et al., 2023)—and does not capture higher-order abilities like in-context reasoning (Wei et al., 2022; Kojima et al., 2022) or adaptive self-refinement (Madaan et al., 2023) – unless explicitly trained for them. In contrast, emergent ICL in LMs generalizes broadly and flexibly across diverse problem types (Srivastava et al., 2023). These behavioral and mechanistic differences support our working assumption that Meta-ICL and emergent ICL are fundamentally distinct phenomena (Shen et al., 2024; Mishra et al., 2025). Throughout, our focus in this work is on the latter – the organic emergence of in-context learning in large pretrained models.
ICL $\neq$ Zero-shot learning: Existing literature in the domain of timeseries studies the zero-shot inference capabilities of next-token predictors, i.e., given a context of a given problem their models can predict the forthcoming values (tokens) (Ansari et al., 2024; Das et al., 2024). This differs from the type of ICL we evaluate, which explicitly relies on few-shot learning of a transformation, not zero-shot adaptation. In fact, in Fig. 2, we report that Evo2’s performance begins far below a naive mode baseline (simply guessing the most common output) at one shot, indicating very poor zero-shot capability. Evo2’s ability only becomes meaningful only when many shots are provided. This makes the type of ICL we demonstrate qualitatively different from zero-shot ability, as it explicitly only appears in few-shot contexts.
@@
Are there any prior results on emergent ICL in genomic models? Most genomic models to date are encoders (Ji et al., 2021; Dalla-torre et al., 2024), which preclude autoregressive generation and the study of ICL in the standard sense. Among the few autoregressive genomic models, most remain small, e.g., GENA-LM (300M) (Fishman et al., 2024) and scGPT (53M) (Cui et al., 2024). The recently-released Evo2 series (Brixi et al., 2025) includes models up to 40B parameters, comparable in scale to modern LLMs. This model family provides the first viable testbed for probing emergent ICL, which we study here. A related effort, HyenaDNA (Nguyen et al., 2023), trains on million-token genomic sequences and reports ICL-like behaviors; however, these involve meta-ICL mechanisms (e.g., soft prompting and instruction fine-tuning; see their Sec. 4.3), rather than organically emergent ICL. Similarly, Bio-xLSTM (Schmidinger et al., 2024) demonstrates ICL, but only after explicitly training on deliberately constructed few-shot demonstrations instead of natural biological data - another example of Meta-ICL. An additional avenue of research demonstrates that genomic models can perform a form of ICL by copying prior tokens in-context for prediction (Kantroo et al., 2025) – akin to how induction heads function (Olsson et al., 2022a). This, however, doesn’t demonstrate any ability to infer latent functions or apply a novel rule in-context.
@@
Are there instances of emergent ICL in other non-linguistic modalities? While transformers have been extended to many modalities beyond language, genuine cases of emergent ICL arising purely from large-scale pretraining remain rare. In neuro-signal modeling, for instance, EEG-GPT (Kim et al., 2024) is trained with explicit in-context demonstrations, making it an instance of Meta-ICL rather than emergent adaptation. Similarly, multimodal models such as Flamingo (Alayrac et al., 2022), BLIP/BLIP-2 (Li et al., 2022; 2023a), and Emu (Sun et al., 2023; 2024) exhibit in-context behavior only because they are explicitly trained to process demonstration patterns – again, meta-learned rather than emergent. Another example of non-linguistic ICL is Chronos 2 – a timeseries model trained to perform in-context learning through cross-batch learning, where the model’s performance can be improved by learning from multiple time series provided in-context. However, Chronos 2 is trained using exclusively synthetic data designed to elicit cross batch learning and ICL, which firmly places it in the Meta-ICL regime (Ansari et al., 2025). To our knowledge, the only claimed instance of emergent ICL outside language is in vision:  Bai et al. (2024) show that transformers pretrained on natural visual sequences can infer analogical and compositional tasks without explicit supervision. The weakness of this result is that their ICL prompts consist of long sequences of contiguous frames (Bai et al., 2024, Fig 10), which effectively reduces the task to next-frame prediction, rather than true ICL that requires inferring compositions of multiple demonstrations.
@@ 3.1 Experimental Desiderata
Desideratum 1: Cross-domain comparability. Our experimental framework requires that each task be performable by both language and genomic models. Thus, every task must be representable in both linguistic and nucleotide alphabets. This constraint excludes existing benchmarks that rely on domain-specific semantics – such as language reasoning datasets (e.g., BIG-Bench, StrategyQA, GSM8K (Srivastava et al., 2023; Geva et al., 2021; Cobbe et al., 2021)) or biological tasks (e.g. variant effect prediction, exon identification  (Brixi et al., 2025)). While one could, in principle, translate domain-specific tasks into an alternate alphabet (e.g., mapping language tokens to base sequences via quaternary encoding), doing so inherently biases the evaluation: the source domain retains advantages, while the target domain must operate on representations that are unnatural to it. As a result, such cross-domain encodings confound the comparison, reflecting representational translation artifacts rather than genuine differences in ICL behavior.
Desideratum 2: Limited vocabulary. Since genomic models operate on only four nucleotides (A, T, C, G), tasks must be expressible within an equally compact alphabet. This rules out the existing symbolic reasoning and analogy benchmarks (Hodel and West, 2023; Lewis and Mitchell, 2025; Webb et al., 2025), which rely on richer vocabularies or background knowledge (e.g., shapes, colors, or linguistic tokens with explicit semantic/lexical roles). While some of these tasks are more symbolic in nature, they still rely on pretrained knowledge or linguistic instructions about numerical or lexical ordering (i.e. ’e’ coming after ’a b c d’) – information that couldn’t effectively be conveyed in-context to a genomic model (Lewis and Mitchell, 2025; Stevenson et al., 2025). Such tasks cannot be faithfully represented in a four-token regime without introducing artifacts or structural loss. Accordingly, our evaluation focuses on tasks that retain abstract reasoning structure while remaining compatible with the low-vocabulary symbolic space shared across linguistic and genomic models.
@@ 3.2 Experimental Setup
Task Formulation: in-context program induction over abstract symbols. Given our desiderata (§3.1), we design symbolic induction tasks that require a small lexicon. Each task requires a model to infer a hidden transformation rule from a few input–output examples and apply it to a new query. We refer to this process as program synthesis in-context. This setup isolates in-context learning ability by removing the influence of pretrained world knowledge. If the function family $f$ and symbol space $S$ are chosen carefully, the evaluation is far from the model’s pretraining distribution. Similar formulations have been used in prior work on compositional reasoning (Brown et al., 2020a), analogical reasoning via Raven’s Progressive Matrices (Raven et al., 1962; Webb et al., 2023), and more recently in ARC-AGI (Chollet, 2019).
@@ 3.3 Model Families and Selection Rationale
For fair cross-domain comparison, we use two representative model families: Qwen3 for human language and Evo2 for genomics. The rationale for this selection is as follows:
  1. (a)
    Parameter scaling: Both families span multiple orders of magnitude in parameter count, enabling systematic scaling analysis of ICL ability. The Qwen3 series ranges from 0.6B to 14B parameters, while Evo2 includes 1B, 7B, and 40B models (Yang et al., 2025; Brixi et al., 2025). This parallel scaling structure facilitates consistent measurement of how ICL performance evolves with model size across linguistic and genomic modalities.
  2. (b)
    Compute matching: The largest models in each family are trained with comparable total compute, offering an opportunity for an approximately compute-matched cross-domain comparison. Using the standard $6ND$ estimate (Kaplan et al., 2020), Qwen3-14B-Base is trained with about $3.2\times 10^{24}$ FLOPs, while Evo2-40B is trained with $2.25\times 10^{24}$ FLOPs (Yang et al., 2025; Brixi et al., 2025), making Evo2 uniquely well-suited for comparison with Qwen3 at scale.
  3. (c)
    Availability of base models: The Qwen3 family releases base (pre-instruction-tuned) models at all scales up to 14B parameters, enabling direct evaluation of the intrinsic inductive reasoning ability of pure next-token predictors, without instruction-tuning artifacts. The Evo2 base models have no additional instruction tuning applied—the only nuance is that Evo2 1B is trained at a context length of 8192 nucleotides, whereas the 7B/40B are extended to a context length of one million.
  4. (d)
    Tokenizer: Qwen3 uses a standard BPE tokenizer with a vocabulary size of 151,669. Evo2 uses a byte-level tokenizer, so individual nucleotides are mapped to individual tokens. Notably, Qwen’s tokenizer maps single digits to single tokens, allowing for parity in our experiments (Yang et al., 2025; Brixi et al., 2025).
  5. (e)
    Context length: Qwen3’s 0.6B and 1.7B models have a context length of 32K. The remaining dense models have a context length of 128K. Evo2’s 1B model has a context length of 8K, and the remaining models have a context length of 1M. As none of our experiments approach these limits, we can use all models in both families without fear of context length as a confound (Yang et al., 2025; Brixi et al., 2025).
  6. (f)
    Training corpora: All Qwen3 models are trained on 36 trillion text tokens covering a wide variety of topics and languages. Evo1 1B is trained on 1 trillion tokens, Evo2 7B is trained on 2.4 trillion tokens, and Evo2 40B is trained on 9.3 trillion tokens. These extensive training corpora ensure that any potential ICL dynamics have thoroughly emerged (Yang et al., 2025; Brixi et al., 2025).
  7. (g)
    Architecture: Qwen3 is based on a conventional Llama-like transformer architecture, while Evo2 uses the StripedHyena2 architecture which intersperses convolutional layers with attention-based ones. While ideally both model families would use the same architecture, there were no vanilla transformers at Evo2’s compute scale (Yang et al., 2025; Brixi et al., 2025).
  8. (h)
    Licensing: All models are released under Apache 2.0 (Yang et al., 2025; Arc Institute, 2025a; c; b), ensuring reproducibility.